OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 20 inches? Use π ≈ 3.14. A. 62.8 inches B. 120 inches C. 125.6 inches D. 251.2 inches I THINK IT IS c< PLEASE HELP< I WILL GIVE MEDAL AND FAN!!!

OpenStudy (anonymous):

@mathmath333 am I right

OpenStudy (anonymous):

i am pretty sure i am right, i just came to see if i am

OpenStudy (anonymous):

i used the circle calculater to help

OpenStudy (anonymous):

nv, i got it

OpenStudy (mathmath333):

\(\large C=2\pi r=2\times 3.14\times 20=125.6\)